OpenStudy (anonymous):

Check my answer?? What is the d value of the following arithmetic sequence? 5, 2, -1, -4, -7,… I got -3

OpenStudy (anonymous):

Well generally when we have a arithmetic sequence we use the following "formula" not really a formula that's why I put it in quotations :P \[\huge t_2 - t_1 = t_3 - t_2 = d~~~~ \text{t = term}\]

OpenStudy (anonymous):

\[2-5=-1-2 \implies - 3\] so you're good :)

OpenStudy (anonymous):

Thank you!!!